OpenStudy (anonymous):

The circumference of a circle is 65π. In terms of pi, what is the area of the circle? A) 600 π B) 936.5 π C) 950.25 π D) 1056.25 π

OpenStudy (jhannybean):

\[C = 2\pi r\] if we solve for \(r\), we get \(r=\dfrac{C}{2\pi}\) Then we look at the area of a circle: \(A = \pi r^2\). All we have to do now is substitute what we have solved for \(r\) \[A=\pi \left(\frac{C}{2\pi}\right)^2\]

OpenStudy (anonymous):

ok im really rusty so i forget the steps